DX LUCI/AO Commissioning

03:10 Finally opened when winds dropped below 20 m/s

03:25 LUCI Script
  • Preset AO to 11. 3 mag: Success
  • AcquireRefAO: no Ref Star: Failed
  • bad coordinates for AO Ref Star. Fixed

03:45 LUCI Script
  • Preset AO to 11. 3 mag: Success
  • AcquireRefAO: Success
    • Optimize gain give 2.5, 2.0, 2.0: Too high
    • Set to 2.25, 1.2, 1.0
  • Open loop (same time on both sides) due to seeing spike

04:10 LUCI Script
  • Preset AO to 11. 3 mag: Success
  • AcquireRefAO: Success
    • Optimize gain give 2.5, 2.0, 2.0: Too high
    • Set to 2.25, 1.2, 1.0

04:43:09 Offset failed because Loop

04:46:49 Loop Open due to safe skip at ~3:00
  • 573, 574, 575, 562, 13, 3 Rolled edge
  • Low gain 1.35, 0.65, 0.45 No problems TN=050304
  • 1.85 0.8 0.65 TN = 050400
  • 1.6 to 1.9 TN = 051319
  • 1.9 to 2.45 TN 051348
  • 2.45 TN 051446
  • Opened loop and set flat to get rid of rolled edge at 3:00
  • 051556 See a bit
  • 051610
  • 051631 not much
  • 051656 not much
  • 051730 not much

05:24 Offset sent from LUCI

05:24 Loop opened with safe skip. Striping at 11:00

05:31 All these problems are at bin 2. These are old reconstructors. Do we need to remeasure bin 2, 3, 4?

05:36 LUCI script with bin 1 reference star

054945, 055208, 055441, 055625

05:58 After offset the star was not on the hotspot. Loop closed and had to pull star in. Large tip and then open loop
  • 055805 during skip frame and loop open

06:02 New preset out of the wind

06:20 Closed for wind

08:26 Looking at telemetry data
  • 20170922_551610
  • From pos[200] looks like it may be actuator 566
  • Put in pos_do_not_use list

-- %USERSIG{DougMiller - 2017-09-22}%
Topic revision: r2 - 22 Sep 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback